40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14093     Entry time: Fri Jul 20 22:53:15 2018
Author: Koji 
Type: Update 
Category: ASS 
Subject: Attempt to resurrect Yarm ASS  

[Koji Gautam]

We managed to realize stable ASS configuration for Yarm. The transmission of 1.06~1.07 was recovered by introducing intentional beam spot offset in the horizontal direction towards the opposite side of the elliptic reflector. The end table optics were adjusted to have the spots about the center of the mirrors, lenses, and PDs/QPDs.


Preparation

- The Y arm was manually aligned with a given input axis. The transmission was ~0.8.
- Then, TT2 was moved in yaw such that it introduced the horizontal beam shift at the end. By moving the spot to the opposite side of the reflector. The transmission ~0.95 was obtained after patient alignment work.

- Went to the end table and checked the spots. The beam was not at the center of the last 1" lens for the Trans PDs. The beam steering was adjusted to have the spot nicely going through the lens and the mirrors. This made the transmission level to be ~1.05.

- The beam centering on the Trans PD was checked and adjusted.
- The beam centering on the RF BBPD for the arm scan was checked. The spot was too big for that PD. The lens was slightly moved away from the PD to make the spot on the BBPD small. Now the PD saw the plateu when the steering was scanned (i.e. the spot is small enough).

- With the Y arm locked with MC2, the servo gain needs to be 0.012 instead of nominal 0.015 with ETMY to prevent from servo oscilating.

ASS tuning

- First of all, only the bottom 4 loops out of total 8 loops were tuned. They are the servos for the beam alignment with regard to the caivty. The linearity and the zero crossings were checked with regard to the reference alignment. All of these 4 showed offsets that causes the servo running away. Don't know the reason of this offset, but it is freq dependent. Therefore the dither freqs were tuned to make the offset zeroed, and tuned the demod phases there. This kept the transmission as high as the reference (~1.05)

- This allowed us to play with the spot position a bit by tuning the caivty alignment. In the end, the transmission of ~1.08 was obtained. Using this alignment, A2L offset for ETMY Yaw was determined to be +17 (to make the error signal -17). This offset produces almost a beam radius (5mm) shifted on the end mirror towards the opposite direction of the reflector.

- The nominal servo setting made the spot servo running away. Gautam pointed out that this could be a gain hierarchy problem (i.e. the spot servos are too fast). We ended up reducing the gain of the servo from 1.0 to 0.3 to make the spot servo stable.

- All the ASS setting was stored in a new snap file "script/ASS/ASS-DITEHR_ON.snap". The previous snap was saved to "script/ASS/ASS_DITHER_ON_preVent201807.snap". This did not save the exc gains of the oscillators. Therefore "DITHER_ASS_ON.py" was modified to have the new exc gains (CLKGAIN). The old values are stored in the comments in this script.


Overall this is not an ideal situation as we don't know what is the actually cause of the offsets in the dither error signals. We expect to correct the beam clipping and the suspension sooner or later. Therefore, we will come back to the ASS again once the other issues are corrected.

Attachment 1: 02.png  80 kB  Uploaded Sat Jul 21 01:43:20 2018  | Hide | Hide all
02.png
ELOG V3.1.3-